Find the value of a and b if

lim x(1+acosx)-bsinx/ (x^3)=1

x→∞


1
Expert's answer
2021-06-16T11:25:08-0400

Solution:-

there is error in question as i got when I solved the full question that is it comes out to be infinite , the limit should tend to zero instead of infinite to get the result , so i am solving the problem by taking limit as tends to zero.


"lim_{x \rightarrow 0} {x(1+acosx)-bsinx\over (x^3)}=1"

as we see it is "{0 \over 0}" type of limit

using L' Hospital rule

we get

="lim_{x \rightarrow 0} {1+acosx-axsinx-bcosx\over (3x^2)}=1"

="lim_{x \rightarrow 0} {1+(a-b)cosx-axsinx\over (3x^2)}=1"

we see that limit exist when a-b=-1 , the "0\over 0" exist with this condition only

To use L'Hospital , we see a-b has to be -1


"\therefore" "a-b =-1" ............................................................(1)

Applying L'Hospital again ,

= "lim_{x \rightarrow 0} {-(a-b)sinx-asinx-axcosx\over (6x)}=1"

this is is in form of "{0 \over 0}" , using L'Hospital yet again ,

= "lim_{x \rightarrow 0} {(b-2a)cosx+axsinx-acosx\over (6)}=1"

="lim_{x \rightarrow 0} {(b-3a)cosx+axsinx\over (6)}=1" { sin(0)=0 , cos(0)=1}

"\implies"

when we put x=0 , we get cosx =1, sinx=0, and putting the limit equals to 1 we get the following result

b-3a=6 ...........................................................................(2)

solving (1) and (2) we get

a="-5\over 2"

b="-3\over 2"